# SplAgger: Split Aggregation for Meta-Reinforcement Learning

### 05 Mar 2024 v1

@@ Abstract
A core ambition of reinforcement learning (RL) is the creation of agents capable of rapid learning in novel tasks. Meta-RL aims to achieve this by directly learning such agents. One category of meta-RL methods, called black box methods, does so by training off-the-shelf sequence models end-to-end. In contrast, another category of methods have been developed that explicitly infer a posterior distribution over the unknown task. These methods generally have distinct objectives and sequence models designed to enable task inference, and so are known as task inference methods. However, recent evidence suggests that task inference objectives are unnecessary in practice. Nonetheless, it remains unclear whether task inference sequence models are beneficial even when task inference objectives are not. In this paper, we present strong evidence that task inference sequence models are still beneficial. In particular, we investigate sequence models with permutation invariant aggregation, which exploit the fact that, due to the Markov property, the task posterior does not depend on the order of data. We empirically confirm the advantage of permutation invariant sequence models without the use of task inference objectives. However, we also find, surprisingly, that there are multiple conditions under which permutation variance remains useful. Therefore, we propose SplAgger, which uses both permutation variant and invariant components to achieve the best of both worlds, outperforming all baselines on continuous control and memory environments.
@@ 1 Introduction
In contrast, a distinct category of research focuses on methods specialized for meta-RL. These methods typically infer an explicit posterior over tasks, given data collected from a new task. In order to infer a posterior over task, these methods generally make use of distinct objectives and distinct sequence models, designed to enable inference of the unknown task. In particular, it is common to use sequence models that are invariant to the order of their inputs, which we refer to as permutation invariant aggregation. It can be shown that the true posterior over tasks does not depend on this order, following directly from the Markov property. Collectively, these methods are referred to as task-inference methods (Beck et al., 2023a).
While many task inference methods have been developed for meta-RL, recent work has shown that black-box methods to be more effective in practice (Ni et al., 2022; Beck et al., 2023b). However, these results focus primarily on demonstrating the superiority of the end-to-end objective used in black-box methods over the task-inference objective, and do not investigate the particular sequence models. This has left an open question: When using an end-to-end objective, is it still worth using permutation invariant sequence models?
In this paper, we answer in the affirmative. We show that permutation invariant sequence models still confer an advantage in a number of domains, even when trained end-to-end. Surprisingly, we still do find domains where dependence on the permutation remains useful. Specifically, we find sequence models with a permutation variant component to be less sensitive to choices in the permutation invariant component, and we find permutation variance useful when there exist permutation variant sub-optimal policies. We extensively investigate the conditions under which each type of sequence model is useful, conduct analysis to support our conclusions, and propose a simple sequence model, called Split Aggregator, or SplAgger, adapted and simplified from the literature on partial observability (Beck et al., 2020). SplAgger, depicted in Figure 0(c), carefully uses both permutation invariant and permutation variant components to achieve the best of both worlds and high returns in all domains evaluated.
@@ 2 Related Work
Task-inference methods explicitly attempt to infer a posterior distribution over the identify of the task (Beck et al., 2023a). Following directly from the Markov property, it can be shown that this posterior does not depend on the order of the data on which the agent conditions. While generic sequence models, such as RNNs, may model permutation invariance, they must learn to do so. In order to incorporate this inductive bias directly, methods generally modify the sequence model to be permutation invariant (Rakelly et al., 2019; Galashov et al., 2019; Raileanu et al., 2020; Wang & van Hoof, 2022; Imagawa et al., 2022). One popular method, called probabilistic embeddings for actor-critic RL, or PEARL (Rakelly et al., 2019), incorporates order-invariance into the probability density function of a stochastic latent variable summarizing history. Specifically, the density function, modelled as a product over individual transitions in the data, is permutation invariant. We compare to this style of aggregation in our experiments. Another approach uses commutative operators applied across the data (Imagawa et al., 2022; Wang & van Hoof, 2022; Galashov et al., 2019). Generally, these can be viewed as (conditional) Neural Processes (CNP) (Garnelo et al., 2018b; a), and so we compare to this style of aggregation as well. Yet another approach uses attention, self-attention, or transformers (Mishra et al., 2018; Fortunato et al., 2019; Nguyen & Grover, 2022). Attention is inherently permutation invariant. Still, attention is computationally expensive: whereas both commutative aggregation and recurrent networks use $O(1)$ memory and compute per timestep, attention generally requires $O(t^{2})$ memory and compute per timestep $t$. While fast approximations of attention exist (Katharopoulos et al., 2020), the computation requirements are significantly larger. Additionally, our limited experimentation shows them to have difficulty learning on our domains. (See the Appendix for more.) Thus, we limit our solutions to constant memory and compute, in line with sequence models designed to quickly handle long contexts (Garnelo et al., 2018b; Beck et al., 2020). Finally, we compare to Aggregated Memory for RL, or AMRL (Beck et al., 2020). While AMRL was originally proposed as a method for POMDPs, it was evaluated in meta-RL problem settings. Our method proposes a simplification to AMRL that is vital in practice. Details of AMRL are covered in Section 4.
@@
The methods we investigate in this paper can be seen as performing in-context learning. Learning that occurs after training and within the activations of a sequence model is called in-context learning (Brown et al., 2020). Black box and task inference methods both perform in-context learning. In part due to the popularity of large language models (Devlin et al., 2019; Brown et al., 2020; Chowdhery et al., 2022), in-context learning has gained significant traction recently, including in application to decision making problems (Raparthy et al., 2023; Lee et al., 2024) and reinforcement learning (RL) (Kirsch et al., 2022). While other meta-RL methods exist that do not use sequence models for in-context learning, such as parameterized policy gradient methods, they generally require more samples to adapt to novel tasks, including those used in our benchmarks (Zintgraf et al., 2021; Beck et al., 2023a). In-context learning promises to address the sample inefficiency still impeding progress in reinforcement learning. Learning to perform in-context RL is the problem studied in meta-RL.
@@ 3.1 Problem Setting
Meta-RL algorithms learn an RL algorithm, $f(\tau)$, over a distribution of MDPs, or tasks. $f(\tau)$ maps from the data, $\tau$, sampled from a single MDP, $\mathcal{M}\sim p(\mathcal{M})$, to policy parameters $\phi$. As in a single RL task, $\tau$ is a sequence up to time-step $t$ forming a trajectory of states, actions, rewards, and next states, $\tau\in(\mathcal{S}\times\mathcal{A}\times\mathbb{R}\times\mathcal{S})^{t}$. Here, $\tau$ may span multiple episodes within a single MDP, since multiple episodes of interaction may be necessary for learning. Collectively, we refer to these episodes as a meta-episode, and use the same symbol, $\tau$, to refer to it. The policy, $\pi(a|s;\phi=f_{\theta}(\tau))$, is parameterized by $\phi$. $f$ is represented as a sequence model, parameterized by $\theta$, which are referred to as the meta-parameters.
@@
3.2 Permutation Invariance
$P(\mathcal{M}|\tau)$ (2)
$\propto P(\tau|\mathcal{M})P(\mathcal{M})$ // Bayes’s rule (3)
where $\tau_{t}$ is shorthand for the transition, $(s_{t},a_{t},r_{t},s_{t+1})$, and the full proof is in the Appendix.
In this expression, the posterior, $P(\mathcal{M}|\tau)$, does not depend on the order of the transitions, $\tau_{1},..,\tau_{t}$. While it is possible to learn each $P(\tau_{i}|\mathcal{M})$, this form is not particularly amenable to inference, since it requires marginalizing over all MDPs at test time. Still, it is possible to incorporate the permutation invariant structure into the sequence model directly.
$g(e_{1},...,e_{t})=e_{1}\bigoplus e_{2}\bigoplus...\bigoplus e_{t}.$ (5)
@@ 3.3 AMRL
Here, the RNN is able to handle short permutation variant sequence, which can then be integrated without respect to order by the permutation invariant aggregation over larger periods of time. Since we split the neurons before aggregation, we call this process split aggregation.
Additionally, AMRL suggests modifying each Jacobian, $\frac{dg}{de_{i,2/2}}\forall_{i}$, when computing the chain rule in the backward pass. Specifically, AMRL overwrites the true Jacobian with the identify matrix, $I$. This is called passing the gradient straight through, or an ST gradient modification. Since the sum aggregator actually has the Identify as the true Jacobian, this can be seen as replacing the gradient with that of a sum. For the average and max aggregators investigates, the Jacobian is already similar to the Identify. Specifically, the average has a Jacobian that is $I/t$, and the max as an expected Jacobian that is $I/t$, under mild assumptions. Thus, the modification is hypothesized can be seen as a re-scaling of the gradient that does not diminish with time. It is hypothesized that the ST modification has no negative impacts while also preventing gradient decay. While the ST modification does prevent one type of gradient decay, the other aggregators already address the relevant type of gradient decay, and the ST aggregator causes both the explosion of other gradients and a severe decrease in performance. We will show that this is the case in Sections 5 and 6, motivating the need for SplAgger.
@@
4 SplAgger
Here we present our model, Split Aggregator, or SplAgger. As motivation, we first present a preview of our later results. In Figure 2, we experiment with the permutation invariant model using a point-wise maximum aggregator, as suggested in AMRL, and a permutation variant model, the RNN. We can see that permutation variance improves performance on the MC-LS domain, but decreases on the Planning Game. Moreover, when we add the ST gradient modification, we can see that performance decreases severely, in Figure 2. These result motivate the need for our model, SplAgger, which achieves the highest returns in both domains.
SplAgger uses the same split aggregation as in AMRL, but without the ST gradient modification. Simplifying AMRL by removing this gradient modification is key to its performance. Additionally, SplAgger uses a different architecture for the policy, compared to the original AMRL paper. Recently, it has been shown that the hypernetwork (Ha et al., 2017) architecture to be critical in unlocking the performance of end-to-end objectives and enabling black box methods to have superior performance to task inference methods (Beck et al., 2023b). Thus, main insight of SplAgger is to add AMRL to hypernetworks trained end-to-end, but to remove the ST gradient modification. SplAgger is the combination of these components, and is depicted in Figure 0(c).
@@ 5 Experiments
On the four primary benchmark environments, we compare to four baselines. Since prior results demonstrate the need for hypernetworks when training end-to-end (Beck et al., 2023b), all baselines have been evaluated using hypernetworks, with design choices detailed in the Appendix. We additionally present negative results on a novel initialization method in the Appendix. The baselines evaluated are listed below.
@@
For the RNN baseline, we use a standard gated recurrent unit (Cho et al., 2014), as in Zintgraf et al. (2021) and Beck et al. (2023b).
PEARL-style Aggregation.
The PEARL algorithm (Rakelly et al., 2019), incorporates order-invariance into the probability density function of a stochastic latent variable summarizing history. The density function is modelled as a product over individual transitions in the data. We compare to the style of aggregation used in this algorithm. Details on the additional design choices and hyperparameters for this model are presented in the Appendix.
@@
We compare to the AMRL aggregation method (Beck et al., 2020), presented in Section 3. We use the point-wise maximum aggregator in our experiments, both to match the aggregator used in SplAgger, and because that aggregator was found to be strongest by Beck et al. (2020).
CNP.
Finally, we compare to the CNP aggregator (Garnelo et al., 2018a). CNP consists of order-invariant aggregation without any additional components. Specifically, $f(\tau)=g(e_{1},...,e_{t})$. Here, we use mean aggregation for $g$, as suggested by Garnelo et al. (2018a).
@@ 5.1 MuJoCo Benchmarks
The first two environments for benchmarking are variants of MuJoCo proposed by Zintgraf et al. (2021), and both involve legged locomotion. While these environments have no great demands on memory, they are common meta-RL benchmarks (Humplik et al., 2019; Rakelly et al., 2019; Zintgraf et al., 2021; Beck et al., 2022; 2023b), which enable us to evaluate what effect SplAgger has on standard RL tasks. In Cheetah-Dir, the agent outputs six different torques in order to control a cheetah robot. The agent is rewarded for its velocity in either the forward or reverse direction, depending on the sampled task, and receives a penalty for the magnitude of the torques. The agent receives a 17-dimensional observation consisting of the position, angle, and velocity of each body part. In Walker, the agent likewise outputs six torques and receives 17-dimensional observations. However, in Walker, the agent controls a two-legged torso morphology, and the agent receives a reward for running only in the forward direction. The tasks are defined by random samples of 65 different physics coefficients, such as body mass and friction, which collectively define the task.
Results are shown in Figure 3. On Cheetah-Dir, while the improvement in return is not large, SplAgger does achieve the greatest return. Training PEARL on this domain is unstable and PEARL receives significantly lower returns. On Walker, we see similar performance across all methods. Overall, SplAgger achieves similar or greater returns compared to other baselines. This demonstrates that our method, designed to improve environments with difficult demands on memory, also does not decrease performance on domains with limited memory requirements.
@@
Figure 4: Depiction of Planning Game, T-LS, T-Maze Agreement, and T-Maze latent environments used in Sections 5.2 and 5.3.

### 08 Mar 2024 v2

@@ Abstract
A core ambition of reinforcement learning (RL) is the creation of agents capable of rapid learning in novel tasks. Meta-RL aims to achieve this by directly learning such agents. Black box methods do so by training off-the-shelf sequence models end-to-end. By contrast, task inference methods explicitly infer a posterior distribution over the unknown task, typically using distinct objectives and sequence models designed to enable task inference. Recent work has shown that task inference methods are not necessary for strong performance. However, it remains unclear whether task inference sequence models are beneficial even when task inference objectives are not. In this paper, we present strong evidence that task inference sequence models are still beneficial. In particular, we investigate sequence models with permutation invariant aggregation, which exploit the fact that, due to the Markov property, the task posterior does not depend on the order of data. We empirically confirm the advantage of permutation invariant sequence models without the use of task inference objectives. However, we also find, surprisingly, that there are multiple conditions under which permutation variance remains useful. Therefore, we propose SplAgger, which uses both permutation variant and invariant components to achieve the best of both worlds, outperforming all baselines on continuous control and memory environments.
@@ 1 Introduction
By contrast, a distinct category of research focuses on methods specialized for meta-RL. These methods typically infer an explicit posterior over tasks, given data collected from a new task. To do so, they generally use distinct objectives and distinct sequence models, designed to enable inference of the unknown task. In particular, it is common to use sequence models that are invariant to the order of their inputs, which we refer to as permutation invariant aggregation. Due to the Markov property, the true posterior over tasks does not depend on this order. Collectively, these methods are referred to as task-inference methods (Beck et al., 2023a).
While many task inference methods have been developed for meta-RL, recent work has shown black-box methods to be more effective in practice (Ni et al., 2022; Beck et al., 2023b). However, these results focus primarily on demonstrating the superiority of the end-to-end objective used in black-box methods over the task-inference objective, and do not investigate the effect of the particular sequence model. This leaves an open question: When using an end-to-end objective, is it still worth using permutation invariant sequence models?
In this paper, we answer in the affirmative. We show that permutation invariant sequence models still confer an advantage in a number of domains, even when trained end-to-end. However, we also find, surprisingly, that there are domains where dependence on the permutation remains useful. Specifically, we find sequence models with a permutation variant component to be less sensitive to choices in the permutation invariant component, and we find permutation variance useful when there exist permutation variant suboptimal policies. We extensively investigate the conditions under which each type of sequence model is useful, conduct analysis to support our conclusions, and propose a simple sequence model, called Split Aggregator, or SplAgger, adapted and simplified from the literature on partial observability (Beck et al., 2020). SplAgger, depicted in Figure 0(c), uses both permutation invariant and permutation variant components to achieve the best of both worlds and high returns in all domains evaluated.
@@ 2 Related Work
Task-inference methods explicitly attempt to infer a posterior distribution over the identify of the task (Beck et al., 2023a). Following directly from the Markov property, it can be shown that this posterior does not depend on the order of the data on which the agent conditions. While generic sequence models, such as RNNs, may model permutation invariance, they must learn to do so. In order to incorporate this inductive bias directly, methods generally modify the sequence model to be permutation invariant (Rakelly et al., 2019; Galashov et al., 2019; Raileanu et al., 2020; Wang & van Hoof, 2022; Imagawa et al., 2022). One popular method, called probabilistic embeddings for actor-critic RL, or PEARL (Rakelly et al., 2019), incorporates permutation invariance into the probability density function of a stochastic latent variable summarizing history. Specifically, the density function, modelled as a product over individual transitions in the data, is permutation invariant. We compare to this style of aggregation in our experiments. Another approach uses commutative operators applied across the data (Imagawa et al., 2022; Wang & van Hoof, 2022; Galashov et al., 2019). Generally, these can be viewed as (conditional) Neural Processes (CNP) (Garnelo et al., 2018b; a), and so we compare to this style of aggregation as well. Yet another approach uses attention, self-attention, or transformers (Mishra et al., 2018; Fortunato et al., 2019; Nguyen & Grover, 2022). Attention is inherently permutation invariant. Still, attention is computationally expensive: whereas both commutative aggregation and recurrent networks use $O(1)$ memory and compute per timestep, attention generally requires $O(t^{2})$ memory and compute per timestep $t$, for autoregressive inference. While fast approximations of attention exist (Katharopoulos et al., 2020), the computational requirements are significantly larger. Additionally, our limited experimentation shows they have difficulty learning on our domains (see Appendix F for details). Thus, we limit our solutions to constant memory and compute, in line with sequence models designed to quickly handle long contexts (Garnelo et al., 2018b; Beck et al., 2020). Finally, we compare to Aggregated Memory for RL, or AMRL (Beck et al., 2020). While AMRL was originally proposed as a method for POMDPs, it was evaluated in meta-RL problem settings. Our method proposes a simplification to AMRL that is vital in practice. Details of AMRL are covered in Section 4.
@@
The methods we investigate in this paper can be seen as performing in-context learning. Learning that occurs after training and within the activations of a sequence model is called in-context learning (Brown et al., 2020). Black box and task inference methods both perform in-context learning. In part due to the popularity of large language models (Devlin et al., 2019; Brown et al., 2020; Chowdhery et al., 2022), in-context learning has gained significant traction recently, including in decision-making applications (Raparthy et al., 2023; Lee et al., 2024) and reinforcement learning (RL) (Kirsch et al., 2022). While other meta-RL methods exist that do not use sequence models for in-context learning, such as parameterized policy gradient methods, they generally require more samples to adapt to novel tasks, including those used in our benchmarks (Zintgraf et al., 2021; Beck et al., 2023a). In-context learning promises to address the sample inefficiency still impeding progress in reinforcement learning. Learning to perform in-context RL is the problem studied in meta-RL.
@@ 3.1 Problem Setting
Meta-RL algorithms learn an RL algorithm, $f(\tau)$, over a distribution of MDPs, or tasks. $f(\tau)$ maps from the data, $\tau$, sampled from a single MDP, $\mathcal{M}\sim p(\mathcal{M})$, to policy parameters $\phi$. As in a single RL task, $\tau$ is a sequence up to time-step $t$ forming a trajectory of states, actions, rewards, and next states, $\tau\in(\mathcal{S}\times\mathcal{A}\times\mathbb{R}\times\mathcal{S})^{t}$. We can see $\tau$ as a trajectory of transitions, $\tau_{0},\tau_{1},...,\tau_{t}$, where $\tau_{t}$ is shorthand for the transition, $(s_{t},a_{t},r_{t},s_{t+1})$. Here, $\tau$ may span multiple episodes within a single MDP, since multiple episodes of interaction may be necessary for learning. Collectively, we refer to these episodes as a meta-episode, and use the same symbol, $\tau$, to refer to it. The policy, $\pi(a|s;\phi=f_{\theta}(\tau))$, is parameterized by $\phi$. $f$ is represented as a sequence model, parameterized by $\theta$, which we refer to as the meta-parameters.
@@
3.2 Permutation Invariance
$P(\mathcal{M}|\tau)$ $\propto P(\tau|\mathcal{M})P(\mathcal{M})$ // Bayes’s rule (2)
with the full proof in Appendix A.
In this expression, the posterior, $P(\mathcal{M}|\tau)$, does not depend on the order of the transitions, $\tau_{1},..,\tau_{t}$. While it is possible to learn each factor, $P(a_{t},r_{t},s_{t+1}|s_{t},\mathcal{M})$, as a function of each individual transition, $\tau_{t}$, this form is not particularly amenable to inference, since it requires marginalizing over all MDPs at test time. Still, it is possible to incorporate the permutation invariant structure into the sequence model directly.
$g(e_{1},...,e_{t})=e_{1}\bigoplus e_{2}\bigoplus...\bigoplus e_{t}.$ (4)
@@ 3.3 AMRL
Here, the RNN is able to handle short permutation variant sequence, which can then be integrated without respect to order by the permutation invariant aggregation over longer periods of time. Since we split the neurons before aggregation, we call this process split aggregation.
Additionally, AMRL modfies each Jacobian, $\frac{dg}{de_{i,2/2}}\forall_{i}$, when computing the chain rule in the backward pass. Specifically, it overwrites the true Jacobian with the identity matrix, $I$. This is called passing the gradient straight through, or an ST gradient modification. Since the sum aggregator actually has $I$ as the true Jacobian, this can be seen as replacing the gradient with that of a sum. For the average and max aggregators, the Jacobian is already similar to $I$. Specifically, the average has a Jacobian that is $I/t$, and the max has an expected Jacobian that is $I/t$, under mild assumptions. Thus, the modification can be seen as a rescaling of the gradient that does not diminish with time. Beck et al. (2020) hypothesize that the ST modification has no negative impacts while also preventing gradient decay.
4 SplAgger
Here we present our model, Split Aggregator, or SplAgger. As motivation, we first present a preview of our experimental results. In Figure 2, we experiment with the permutation invariant model using a point-wise maximum aggregator, as suggested in AMRL, and a permutation variant model, the RNN. Permutation variance improves performance on the some domains (Figure 1(a)), but decreases performance on others (Figure 1(b)). Moreover, when we add the ST gradient modification, performance decreases severely. These result motivate the need for our model, SplAgger, which achieves the highest returns in both domains.
SplAgger uses the same split aggregation as in AMRL, but without the ST gradient modification. Simplifying AMRL by removing this gradient modification is key to its performance. While the ST modification does prevent one type of gradient decay, permutation invariant aggregators already address the relevant type of gradient decay. Moreover, the ST aggregator causes both the explosion of other gradients and a severe decrease in performance. We show that this is the case in Sections 5 and 6, motivating the need for SplAgger. Additionally, SplAgger uses a different architecture for the policy than in AMRL. Recent results show that the hypernetwork (Ha et al., 2017) architecture is critical in unlocking the performance of end-to-end objectives and enabling black box methods to outperform task inference methods (Beck et al., 2023b). Thus, the main idea behind SplAgger is to add AMRL to hypernetworks trained end-to-end but remove the ST gradient modification. SplAgger is the combination of these components and is depicted in Figure 0(c).
@@ 5 Experiments
On the four primary benchmark environments, we compare to four baselines. Since prior results demonstrate the need for hypernetworks when training end-to-end (Beck et al., 2023b), all baselines have been evaluated using hypernetworks, with design choices detailed in Appendix B. We additionally present negative results on a novel initialization method in Appendix E. The baselines evaluated primarily differ in their choice of aggregation function, $g$, and encoding of inputs, $e_{t}$. The baselines are described below.
@@
The RNN baseline can be written $f(\tau)=e_{t}=RNN(\tau_{1},...,\tau_{t})$. Here there is no aggregation function, $g$, and the baseline uses a standard gated recurrent unit (Cho et al., 2014), as in Zintgraf et al. (2021) and Beck et al. (2023b).
CNP.
The conditional neural process (CNP) consists of permutation invariant aggregation without any additional components (Garnelo et al., 2018a). Specifically, $f(\tau)=g(e_{1},...,e_{t})$. Here, $e_{t}$ is a linear encoding of $\tau_{t}$. We use the mean operator for $g$, as suggested by Garnelo et al. (2018a).
@@
AMRL (Beck et al., 2020) uses an RNN to encode $e_{t}$ in addition to permutation invariant aggregation, and is described in Section 3. For AMRL, we use the pointwise maximum aggregator in our experiments, both to match the aggregator used in SplAgger, and because that aggregator was found to be strongest by Beck et al. (2020).
PEARL.
The PEARL baseline uses the aggregation method from the PEARL algorithm (Rakelly et al., 2019), which incorporates permutation invariance into the probability density function of a stochastic latent variable summarizing history. The density function is modelled as a product over individual transitions in the data. We can write this as $f(\tau)=g(e_{1},...,e_{t})=z\sim\alpha\prod_{t=1}^{t=T}\mathcal{N}(z;\mu_{t}=e_{t,1/2},\sigma^{2}_{t}=diag(e_{t,2/2}))$, where $e_{t}$ is a linear encoding of $\tau_{t}$ and $\alpha$ is a normalizing constant. To compare the effects of aggregation in isolation, our PEARL baseline only implements the aggregation method used in PEARL, and leaves the rest of the algorithmic choices the same as in SplAgger. Additional design choices and hyperparameters for PEARL are presented in Appendices B and C.
@@ 5.1 MuJoCo Benchmarks
The first two environments for benchmarking are variants of MuJoCo proposed by Zintgraf et al. (2021), and both involve legged locomotion. While these environments have no great demands on memory, they are common meta-RL benchmarks (Humplik et al., 2019; Rakelly et al., 2019; Zintgraf et al., 2021; Beck et al., 2022; 2023b) that enable us to evaluate what effect SplAgger has on standard RL tasks. See Appendix H for details on the environments.
Results are shown in Figure 4. SplAgger achieves the greatest return, though the improvement is modest. Training the PEARL baseline on this domain is unstable and PEARL receives significantly lower returns. On Walker, we see similar performance across all methods. Overall, SplAgger achieves similar or greater returns compared to other baselines. This demonstrates that our method, designed to improve environments with difficult demands on memory, also does not decrease performance on domains with limited memory requirements.
@@
Figure 3: Depictions of the Planning Game, T-LS, T-Maze Agreement, and T-Maze Latent environments used in Sections 5.2 and 5.3.